can you show me step by step how to solve the following equation; (x+1)/(x^2+2x)-(x+4)/(x^2+x)=(-3)/(x^2+3x+2)
Answer by ankor@dixie-net.com(22740)


Factor the denominators

Multiply by common denominator x(x+1)(x+2), Cancel the denominaors, results:
(x+1)(x+1) - (x+2)(x+4) = -3x
FOIL
x^2 + 2x +1 - (x^2 + 6x + 8) = -3x
remove the brackets change the signs
x^2 + 2x +1 - x^2 - 6x - 8 = -3x
Combine like terms
x^2 - x^2 + 2x - 6x + 3x + 1 - 8 = 0
-x - 7 = 0
-x = +7
x = -7
:
:
You can confirms this by replacing x with -7 in the original problems